Circular No. 7573

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 B2 (NEAT) E. F. Helin, S. Pravdo, and K. Lawrence, Jet Propulsion Laboratory,

report that their CCD images of this comet taken on Jan. 24.6 and 25.6 UT with the 1.2-m NEAT telescope at Haleakala show a coma diameter of ∼ 29′′; there is no convincing evidence for a tail, though some images suggest a certain asymmetry toward the northwest. Additional astrometry is reported on MPEC 2001-B47, together with the following preliminary parabolic orbital elements by B. G. Marsden:

T = 2001 June 6.094 TT  ω = 336.^{o}966
Ω = 145.566 }2000.0
q = 5.21621 AU  i = 150.249
2001 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Jan. 21 12^{h}07.^{m}71 −22^{o}18.^{′}6 4.916 5.316 108.^{o}8 10.^{o}1 14.7
26 12 03.80 −22 00.2 4.825 5.309 114.5  9.7 14.7
31 11 59.47 −21 37.9 4.737 5.302 120.2  9.2 14.6
Feb. 5 11 54.73 −21 11.4 4.656 5.295 126.0  8.7 14.6
10 11 49.61 −20 40.5 4.580 5.289 131.8  8.0 14.5
15 11 44.15 −20 05.0 4.512 5.283 137.5  7.3 14.5
20 11 38.39 −19 25.0 4.452 5.277 143.1  6.5 14.5
25 11 32.39 −18 40.3 4.402 5.271 148.5  5.6 14.4
Mar. 2 11 26.22 −17 51.2 4.361 5.266 153.5  4.8 14.4

COMETS C/2001 A4 AND C/2001 B3 (SOHO) Further to IAUC 7572, D. Hammer reports his measurements for two

additional Kreutz sungrazing comets, found in C3 coronagraph data at the SOHO website by M. Boschat (C/2001 A4) and M. Oates (C/2001 B3). Though C/2001 B3 was visible in many C3 images, it faded before it entered the field-of-view of the C2 coronagraph. The reduced observations and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 A4 Jan. 14.138  19^{h}59.^{m}7  −23^{o}36′  2001-B21
C/2001 B3  24.388  20 49.7  −21 06  2001-B45
2001 January 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green